Spot FX rate is 0.3500 USD/PLN. You can enter a 2-year forward contract with an exercise price of 4.0000 PLN/USD. market and 10% for USA market, what is the theoretical price of this contract? Is arbitrage possible-if yes, explain why? If arbitrage is possible, what strategy should investor choose in order to get a riskless profit? The rates are fixed in the whole investment period and equal 5% p.a. for Polish
Answer #1
Theoretical price of the contract per IRPT= Spot rate*(1+iUS)/(1+iPLN) = 0.3500*(1.10/1.05)^2 = 0.3841
The forward rate USD/PLN = 1/4 = 0.2500
As the forward price is different, there is possibility for
Covered Interest Rate Arbitrage.
The premium/(discount) in the forward rate = (0.2500/0.3500)^(1/2)-1 = -15.48%
Interest rate differential = 10%-5% = 2.50%
As the interest rate differential is less than the forward discount,
it would be profitable to borrow in the currency having higher
interest rate and then to invest the money in the currency having
lower interest rate.
STEPS TO BE TAKEN ON DAY 1:
*Borrow in $ at 10%
*Convert the $ into PLN at the spot rate of 0.3500$/PLN
*Invest the PLN at 5%
*Enter into a forward contract to sell PLN on the due date of the
deposit.
STEPS TO BE TAKEN AT EOY 2:
*Realize the PLN deposit, with interest at 5%
*Convert the PLN so received into $
*Pay the $ loan with 10% interest, with the dollars received above
and keep the balance $, which represents profit from CIA.